OpenStudy (anonymous):

(1+ i) (-1+2i) simplify

OpenStudy (anonymous):

-1+2i-i+2i^2 when multiplied. i^2=-1 so -1+2i-i-2= -3+i

OpenStudy (anonymous):

-1+2i-i+2i^2 -1+i+2i^2 -1+i+2(-1) -3+i

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

use the FOIL method for factoring. first term in first parenthesis x first term in second parenthesis. Then add that to: first term in first parenthesis x 2nd term in second parenthesis. Add that to: second term in first parenthesis x first term in second parenthesis. Then add that to: second term in first parenthesis x second term in second parenthesis. I know that sounds confusing so here is a picture:

OpenStudy (anonymous):